A hiker walks a certain distance. If he had gone 1 km/h faster, he would have walked the distance in 4/5 of the time. If he had walked 1 km/h slower, he would have taken 2.5 hours longer to travel the distance. Find the distance.


1
Expert's answer
2021-02-10T15:15:59-0500

Let the time taken by the hiker be t hrs.

and Distance be x km and speed of hiker be y km/h

then Distance =speed "\times" time

"\Rightarrow x=y\times t~~~~~-(1)"

Case1:- when hiker travel 1 km/h faster, Time taken="\dfrac{4}{5}" t hrs.

then Distance=speed"\times" time

"\Rightarrow x=(y+1)(\dfrac{4}{5}t)~~~~~~-(2)"


Case 2:- When hiker travels 1 km/h slower, time taken= t+2.5

then "x=(y-1)(t+2.5)~~~~~-(3)"


Substitute eqs.(1) in eqs.(2) and eqs.(3)

"\Rightarrow yt=(y+1)(0.8t)
\Rightarrow y=0.8y+0.8\\\Rightarrow y-0.8y=0.8\Rightarrow 0.2y=0.8\Rightarrow y=4" km/h


Substituting the value of y in eqs.(1) and (3) and solving -

"x=4t,x=(4-1)(t+2.5)"


"\Rightarrow 4t=3(t+2.5)\Rightarrow 4t=3t+7.5\Rightarrow t=7
.5hrs"


Hence The distance "x=4t=4\times 7.5=30km"